# Temporal Preference Optimization for Long-Form Video Understanding

### v2

@@ 4.2 Results
  • Hound-DPO [66] is a previous method that applies Direct Preference Optimization (DPO) [44] on video-LMMs. Their approach leverages ChatGPT [2] to generate ratings for preference data, resulting in a dataset of 17k samples. In contrast, TPO employs a smaller preference dataset generated through a self-generation pipeline, offering a more streamlined alternative.
The primary experimental results, presented in Table 1, compare TPO against the baseline methods on LongVA. The results consistently indicate that LongVA-TPO achieves superior performance, with improvements of 2.9%, 3.1%, and 2.5% on LongVideoBench [55], MLVU [70], and Video-MME [15], respectively. These findings underscore TPO’s capacity to enhance the long-form video understanding capabilities of a pre-trained video-LMM.
@@
In addition, the comparisons between TPO and current state-of-the-art video-LMMs are presented in Table 2. With the introduction of TPO, both the LongVA-TPO and LLaVA-Video-TPO model significantly outperform their corresponding baselines 2.5% and 2.3% on the video-MME benchmark, demonstrating the efficacy of our TPO pipeline. After TPO on LLaVA-Video-7B, our LLaVA-Video-TPO model outperforms all 16 baseline models in the table, including the concurrent work, NVILA [37], as well as several 14B and 20B models, achieving state-of-the-art results on long-form video understanding. The original LongVA model performed worse than Video-CCAM [14] and LongLLaVA [61] on the Video-MME benchmark. However, after incorporating TPO, it successfully outperformed these competitive baselines on Video-MME. Overall, LLaVA-Video-TPO achieves the strongest 7B model on the Video-MME benchmark, setting a new state-of-the-art performance on video comprehension.
Figure 2: The performance of LongVA-TPO and LongVA on MLVU with different input lengths. LongVA-TPO consistently shows performance improvements with longer inputs, whereas LongVA experiences performance degradation when the input exceeds 64 frames.
@@ 4.3.1 Performance with Different Input Frame Count
We evaluate the performance of both our LongVA-TPO model and the original LongVA model across varying input lengths, ranging from 16 to 128 frames, as shown in Figure 2. The results indicate that the LongVA model experiences performance degradation with 128 frames compared to 64 frames. In contrast, our LongVA-TPO model consistently benefits from longer input sequences, leveraging the additional information effectively. This demonstrates the LongVA-TPO model’s robustness in handling extended inputs and its capacity to localize relevant information within long sequences, further validating the efficacy of our TPO approach.
Figure 3: Comparison of our LongVA-TPO model and the original LongVA model on the needle-in-a-haystack task.
@@ 4.3.2 Effect of Dataset Sizes
4.3.3 Effect of different data granularity
In TPO, we design two levels of data granularity to model temporal information at different scales, termed localized TPO and comprehensive TPO. Localized TPO captures local temporal information, while comprehensive TPO models comprehensive, overarching temporal information. To evaluate the individual capabilities, limitations, and necessity of combining localized and comprehensive TPO, we conducted an ablation study. In this study, we maintained the same overall dataset size as the full TPO and trained two separate models based on the LongVA model: one using only Localized TPO data and the other using only Comprehensive TPO data. The performance comparison between these two models and the full TPO model is shown in Table 4.
The performance of localized TPO and comprehensive TPO on the LongVideoBench [55] and VideoMME [15] datasets compellingly demonstrates their effectiveness. However, their limited improvements on the MLVU [70] dataset highlight inherent constraints. When combined, the two methods not only achieve further gains on the LongVideoBench and VideoMME datasets but also deliver a substantial 3% improvement on the MLVU dataset, leading to notable advancements. These findings strongly support the complementary nature of localized and comprehensive TPO, emphasizing the critical importance of constructing multi-granularity preference datasets to fully exploit temporal modeling capabilities.
@@ 4.3.3 Effect of different data granularity
[…]
Table 4: Ablation study of different data granularities.
@@ 4.3.4 Needle-in-a-Haystack
Figure 4: Qualitative comparison between LongVA-TPO model and LongVA on two videos from VideoMME benchmark. Our LongVA-TPO model demonstrates superior generation quality on both comprehensive and localized questions.
The Needle-in-a-Haystack task refers to the challenge of identifying a rare or specific event within a large volume of unstructured video data. Building on the work of Zhang et al. [65], we frame the task using image-based question answering (QA), where images are embedded within a 3-hour-long video, and the model is tasked with answering the corresponding image QA questions. In our experiments, we adopt the same five image QAs as Zhang et al. [65], and present the results in Figure 3. While LongVA, optimized for long-context processing, significantly outperforms LLaVA-NeXT-Video [67] on the Needle-in-a-Haystack task (refer to Figure 4 in [65]), our LongVA-TPO model still demonstrates superior performance, achieving even better results in long-context retrieval.
@@ 4.4 Qualitative Analysis
The qualitative analysis of our LongVA-TPO model and the LongVA model on two videos from the Video-MME benchmark is provided in Figure 4. In the first example, which involves a temporal localization and OCR task, our LongVA-TPO model demonstrates superior performance by accurately localizing the relevant information within the video and providing the correct answer to the OCR question. In the second example, a video discussing the Moon’s formation, LongVA misinterprets the video content by relating it to the Earth’s formation. In contrast, our LongVA-TPO model successfully comprehends and captures the key details of the video’s content.
@@ 5 Related Work
Recently, significant efforts have been devoted to extending the capabilities of large language models (LLMs) [2, 45] into the visual domain, developing various video large multimodal models (video-LMMs), including both proprietary [2, 45] and open-source models [52, 34, 26, 47, 31, 12, 59, 16, 1, 23]. Early approaches focused on curating high-quality video-text instruction-tuning datasets [7, 68, 33, 42, 67], to equip LLMs with video comprehension capabilities. However, these datasets often rely on synthetic data derived from video captions, which limits their effectiveness in capturing visual-temporal dynamics. Other studies have focused on extending pretrained video-LMMs to handle longer video contexts [65, 36, 61, 35, 37, 48], while multimodal interleaved datasets [27, 32] and mixed training strategies [74, 25] have been explored to enhance video-LMM performance. Despite these advancements, the post-training stage for video-LMMs remains underexplored. Recent efforts like LLaVA-Hound [66] utilize ChatGPT to rank model outputs and create preference datasets but fall short in leveraging the temporal information inherent in video data. In contrast, our work pioneers post-training strategies that explicitly incorporate temporal priors to address these limitations.
@@
Temporal grounding is crucial for comprehending the video modality, particularly in long-form videos. Various efforts have been made to enhance temporal localization, including dense captioning [53, 60, 58], highlight detection [24, 40], and temporal video grounding [17, 62, 56], among others. Recent advancements have introduced temporal-aware designs in video large multimodal models (video-LMMs) [46, 9, 29, 21, 51] and have explored the development of agentic systems with temporal grounding capabilities [54]. Unlike these existing approaches, our work focuses on temporal preference optimization during the post-training stage, offering a complementary enhancement to current methods.
@@
Proximal Policy Optimization (PPO) [41, 49, 72] and Direct Preference Optimization (DPO) [44] are two widely used implementations of Reinforcement Learning from Human Feedback (RLHF) [41, 72], serving as key algorithms in preference learning and post-training. In the image-LMM domain, Sun et al. [50] enhanced model visual capabilities by incorporating image captions into the reward modeling process within RLHF. Similarly, Ahn et al. [4] fine-tuned multimodal foundation models using Reinforcement Learning from AI Feedback (RLAIF). Other approaches, such as those proposed by Li et al. [30] and Gunjal et al. [18], directly distilled GPT-4V’s preferences from sampled model responses. A notable strategy involves using text as an intermediate modality, leveraging captions and other descriptive information to extract and distill large language models’ (LLMs) preferences for both images [69] and videos [66]. Furthermore, Pi et al. [43], Zhou et al. [71], and Deng et al. [13] advanced preference learning in image-LMMs by curating preference data through image input manipulation.
@@ 6 Conclusion
We introduced Temporal Preference Optimization (TPO), a scalable post-training framework that enhances temporal grounding in video-LMMs. By leveraging preference learning at both localized and comprehensive levels, TPO effectively captures the intricate temporal dependencies required for long-form video understanding. Extensive experiments across three challenging benchmarks—LongVideoBench, MLVU, and Video-MME—demonstrated TPO’s robust improvements, achieving state-of-the-art performance. By integrating multi-granularity temporal preferences, TPO addresses diverse challenges in long-form video understanding, offering a robust and efficient solution for advancing temporal reasoning in multimodal tasks. One future direction is scaling the preference data to improve coverage and diversity, thereby potentially enhancing TPO’s generalizability. Additionally, while this work focuses on LongVA-7B and LLaVA-Video-7B as representative Video-LMMs, applying TPO to a broader range and larger scale of Video-LMMs would provide insights into its adaptability and performance across different architectures.
Acknowledgements
This work was supported by the National Science Foundation under Grant No. 2026498. Serena Yeung-Levy is a Chan Zuckerberg Biohub — San Francisco Investigator.
Supplementary Material
@@ Appendix B Appendix overview
  • § F Additional Ablation Study on Data Mixture.
  • § G More Qualitative Examples.
@@ Appendix C Implementation Details
We conduct Temporal Preference Optimization (TPO) on LongVA [65] and LLaVA-Video [68], two state-of-the-art video-LMMs. The two TPO models are trained using 8 Nvidia A100 80GB GPUs, with a batch size of 1. For preference optimization, we set the KL-divergence weight ($\beta$) to 0.3 and the supervised fine-tuning (SFT) loss weight ($\alpha$) to 0.5 for LongVA-TPO and we set the KL-divergence weight ($\beta$) to 0.2 and the supervised fine-tuning (SFT) loss weight ($\alpha$) to 1 for LLaVA-Video-TPO. We employ full fine-tuning for both the multimodal projector and the language model while keeping the visual encoder frozen, using a learning rate of $4\times 10^{-6}$ for LongVA-TPO and $3\times 10^{-7}$ for LLaVA-Video-TPO. The training is performed on a curated dataset of 10k samples for one epoch for LongVA-TPO and 10k samples for one epoch for LLaVA-Video-TPO. A cosine learning rate scheduler with a warm-up ratio of 0.1 is utilized [38]. The entire TPO fine-tuning process takes approximately 4 hours on both two models.
For evaluation, we adopt the protocol outlined by LongVA [65] and LLaVA-Video [68], leveraging the official lmms-eval repository [64] to assess our model’s performance on three benchmarks. For LongVA-TPO, we set the parameter $max\_frames\_num=128$ across all three benchmarks. For LLaVA-Video-TPO, we set the parameter $max\_frames\_num=96$ for the Video-MME benchmark and $max\_frames\_num=128$ for the rest of the benchmarks.
@@ Appendix D Preference Dataset Curation
Localized/Comprehensive TPO Ratio LongVideoBench MLVU VideoMME
10:0 53.5 58.7 54.0
8:2 53.8 59.9 54.0
5:5 (Our final TPO model) 54.2 61.7 55.1
2:8 53.4 59.1 54.2
0:10 53.4 58.5 53.8
Table 5: Results of TPO trained on different data mix ratios for localized TPO and comprehensive TPO data based on the LongVA model.
We manually curated a set of 200 keywords assisted with GPT-4o-mini [2], which were utilized to retrieve 8,000 videos from the internet, forming a diverse and comprehensive dataset. Using this dataset, we further developed 10,000 queries paired with their corresponding preference responses, covering a broad range of tasks. The detailed prompts for preference dataset curation are provided in Figure 7 and Figure 8. For LLaVA-Video, we sampled a subset of 10k QA pairs from the LLaVA-Video-178k dataset for comprehensive TPO.
The distribution of video lengths in our collected dataset is presented in Figure 6. The distribution of tasks is illustrated in Figure 6, encompassing Temporal Reasoning (8.7%), Action Reasoning (12.4%), Causal Reasoning (11.1%), Information Extraction (18.0%), Descriptive Questions (12.8%), Summarization (7.5%), Object Reasoning (14.9%), and Spatial Reasoning (13.5%).
@@ Appendix E Preference Dataset Examples
We provide three additional examples of preference datasets, as illustrated in Figure 9. For instance, in Example (a), the task involves an OCR-based query aimed at retrieving the quote located beneath a mural. The dis-preferred response incorrectly identifies the relevant frame, failing to locate the quote below the mural and instead referencing another frame containing the phrase “Forward, Warrior.” In contrast, the preferred response accurately identifies the corresponding frame based on the question. This is achieved by leveraging the highly relevant sub-video segment provided to the video-LMM, enabling the correct extraction of both the quote and its attribution.
@@
Appendix F Additional Ablation Study
Different Data Mix Ratio
We assessed the performance of our TPO model under various data mixing ratios for localized TPO and comprehensive TPO data based on the LongVA [65] model, keeping the total data size constant at 10k. The evaluated ratios included 10:0, 8:2, 5:5, 2:8, and 0:10. The experimental results, summarized in Table 5, clearly demonstrate that the model achieves optimal performance on general video understanding tasks when an equal proportion of localized TPO and comprehensive TPO data is used. This balanced data distribution effectively integrates fine-grained local temporal information with broader high-level temporal context, leading to superior overall model performance.
@@ Appendix F Additional Ablation Study
Appendix G Qualitative Analysis Examples

### v3

@@ 4.2 Results
  • Hound-DPO [66] is a previous method that applies Direct Preference Optimization (DPO) [44] on video-LMMs. Their approach leverages ChatGPT [2] to generate ratings for preference data, resulting in a dataset of 17k samples. In contrast, TPO employs a smaller preference dataset generated through a self-generation pipeline, offering a more streamlined alternative. Besides, to ablate the data source’s effect, we also implement Hound-DPO based on our collected dataset with the same data scale.
The primary experimental results, presented in Table 2, compare TPO against the baseline methods on LongVA. The results consistently indicate that LongVA-TPO achieves superior performance, with improvements of 2.9%, 3.1%, and 2.5% on LongVideoBench [55], MLVU [70], and Video-MME [14], respectively. These findings underscore TPO’s capacity to enhance the general video understanding capabilities of a pre-trained video-LMM.
@@
In addition, the comparisons between TPO and current state-of-the-art video-LMMs are presented in Table 1. With the introduction of TPO, both the LongVA-TPO and LLaVA-Video-TPO models significantly outperform their corresponding baselines 2.5% and 2.3% on the video-MME benchmark, demonstrating the efficacy of our TPO pipeline. After TPO on LLaVA-Video-7B, our LLaVA-Video-TPO model outperforms all 16 baseline models in the table, including the concurrent work, NVILA [36], as well as several 14B and 20B models, achieving state-of-the-art results on video understanding. The original LongVA model performed worse than Video-CCAM [13] and LongLLaVA [61] on the Video-MME benchmark. However, after incorporating TPO, it successfully outperformed these competitive baselines on Video-MME. Overall, LLaVA-Video-TPO achieves the strongest 7B model on Video-MME, setting a new state-of-the-art performance on video comprehension.
Figure 3: Qualitative comparison between LongVA-TPO model and LongVA on two videos from VideoMME benchmark.
@@ 4.3.1 Performance with Different Input Frame Count
We evaluate the performance of both our LongVA-TPO model and the original LongVA model across varying input lengths, ranging from 16 to 128 frames, as shown in Fig. 5. The results indicate that the LongVA model experiences performance degradation with 128 frames compared to 64 frames. In contrast, our LongVA-TPO model consistently benefits from longer input sequences, leveraging the additional information effectively. This demonstrates the LongVA-TPO model’s robustness in handling extended inputs and its capacity to localize relevant information within long sequences, further validating the efficacy of TPO.
Figure 4: Performance comparison of LongVA and LongVA-TPO on the needle-in-a-haystack task across varying input video lengths (horizontal axis) and temporal depths (vertical axis). Heatmaps indicate improved temporal grounding capability of LongVA-TPO.
@@ 4.3.2 Effect of Dataset Sizes
4.3.3 Effect of Post-Filtering
As a critical component of the TPO framework, post-filtering effectively reduces noise and enhances data quality. To further assess its impact, we conducted experiments comparing the performance of LongVA-TPO with and without post-filtering. The results, presented in Table 4, demonstrate that post-filtering consistently improves performance across multiple benchmarks.
@@ 4.3.3 Effect of Post-Filtering
Table 4: Results of LongVA-TPO (TPO) with and without post-filtering. Post-filtering consistently improves performance across multiple benchmarks.
Ratio LongVideoBench MLVU VideoMME
10:0 53.5 58.7 54.0
8:2 53.8 59.9 54.0
5:5 (final model) 54.2 61.7 55.1
2:8 53.4 59.1 54.2
0:10 53.4 58.5 53.8
Table 5: Performance of TPO across different training data mix ratios, varying the proportion of negative responses generated from incomplete versus irrelevant video segments.
4.3.4 Effect of Different Data Mix Ratio
In TPO, we design two different kinds of manipulated schema for the dis-preferred response generation, including creating incomplete and irrelevant videos. To evaluate the individual capabilities, limitations, and necessity of combining incomplete and irrelevant videos, we conducted an ablation study. In this study, we maintained the same overall dataset size as the full TPO and assessed the performance of our TPO model under various data mixing ratios between the negative responses generated from incomplete and irrelevant videos. The evaluated ratios included 10:0, 8:2, 5:5, 2:8, and 0:10. The experimental results, summarized in Table 5, clearly demonstrate that the model achieves optimal performance on general video understanding tasks when an equal proportion of negative responses generated from incomplete and irrelevant videos. This balanced data distribution effectively integrates different type of temporal information, leading to superior overall model performance.
Figure 5: Performance comparison between LongVA-TPO and LongVA on MLVU across varying input lengths. LongVA-TPO consistently benefits from increased input length, whereas LongVA’s performance declines when inputs exceed 64 frames.
@@ 4.3.5 Needle-in-a-Haystack
The Needle-in-a-Haystack task refers to the challenge of identifying a rare or specific event within a large volume of unstructured video data. Building on the work of Zhang et al. [65], we frame the task using image-based question answering (QA), where images are embedded within a 3-hour-long video, and the model is tasked with answering the corresponding image QA questions. In our experiments, we adopt the same five image QAs as Zhang et al. [65], and present the results in Fig. 4. While LongVA, optimized for long-context processing, significantly outperforms LLaVA-NeXT-Video [67] on the Needle-in-a-Haystack task (refer to Fig. 4 in [65]), our LongVA-TPO model still demonstrates superior performance, achieving even better results in long-context temporal localization.
@@ 4.4 Qualitative Analysis
The qualitative analysis of our LongVA-TPO model and the LongVA model on two videos from the Video-MME benchmark is provided in Fig. 3. In the first example, which involves a temporal localization and OCR task, our LongVA-TPO model demonstrates superior performance by accurately localizing the relevant information within the video and providing the correct answer to the OCR question. In the second example, a video discussing the Moon’s formation, LongVA misinterprets the video content by relating it to the Earth’s formation. In contrast, our LongVA-TPO model successfully comprehends and captures the key details of the video’s content.
@@ 5 Related Work
Recently, significant efforts have been devoted to extending the capabilities of large language models (LLMs) [2, 45] into the visual domain, developing various video large multimodal models (video-LMMs), including both proprietary [2, 45] and open-source models [52, 34, 26, 47, 31, 10, 59, 15, 1, 23]. Early approaches focused on curating high-quality video-text instruction-tuning datasets [6, 68, 33, 42, 67], to equip LLMs with video comprehension capabilities. However, these datasets often rely on synthetic data derived from video captions, which limits their effectiveness in capturing visual-temporal dynamics. Other studies have focused on extending pretrained video-LMMs to handle longer video contexts [65, 37, 61, 35, 36, 48, 22], while multimodal interleaved datasets [27, 32] and mixed training strategies [74, 25] have been explored to enhance video-LMM performance. Despite these advancements, the post-training stage for video-LMMs remains underexplored. Recent efforts like LLaVA-Hound [66] utilize ChatGPT to rank model outputs and create preference datasets but fall short in leveraging the temporal information inherent in video data. In contrast, our work pioneers post-training strategies that explicitly incorporate temporal priors to address these limitations.
@@
Temporal grounding is crucial for comprehending the video modality, particularly in long-form videos. Various efforts have been made to enhance temporal localization, including dense captioning [53, 60, 58], highlight detection [24, 40], and temporal video grounding [16, 62, 56], among others. Recent advancements have introduced temporal-aware designs in video-LMMs [46, 7, 29, 20, 51] and have explored the development of agentic systems with temporal grounding capabilities [54]. Unlike these existing approaches, our work focuses on temporal preference optimization during the post-training stage, offering a complementary enhancement to current methods.
@@
Proximal Policy Optimization (PPO) [41, 49, 72] and Direct Preference Optimization (DPO) [44] are two widely used implementations of Reinforcement Learning from Human Feedback (RLHF) [41, 72], serving as key algorithms in preference learning and post-training. In the image-LMM domain, Sun et al. [50] enhanced model visual capabilities by incorporating image captions into the reward modeling process within RLHF. Similarly, Ahn et al. [4] fine-tuned multimodal foundation models using Reinforcement Learning from AI Feedback (RLAIF). Other approaches, such as those proposed by Li et al. [30] and Gunjal et al. [17], directly distilled GPT-4V’s preferences from sampled model responses. A notable strategy involves using text as an intermediate modality, leveraging captions and other descriptive information to extract and distill LLMs preferences for both images [69] and videos [66]. Furthermore, Pi et al. [43], Zhou et al. [71], and Deng et al. [12] advanced preference learning in image-LMMs by curating preference data through image input manipulation.
@@ 6 Conclusion
We introduced Temporal Preference Optimization (TPO), a scalable post-training framework that enhances temporal grounding in video-LMMs. By contrasting between the preference responses from the well-grounded and manipulated video clips, TPO effectively captures the intricate temporal dependencies required for video understanding. Extensive experiments across three challenging benchmarks—LongVideoBench, MLVU, and Video-MME—demonstrated TPO’s robust improvements, achieving state-of-the-art performance. By integrating multi-granularity temporal preferences, TPO offers a robust and efficient solution for advancing temporal reasoning in multimodal tasks. One future direction is scaling the preference data to improve coverage and diversity, thereby enhancing TPO’s generalizability. Additionally, while this work focuses on LongVA-7B and LLaVA-Video-7B as representative Video-LMMs, applying TPO to a broader range and larger scale of Video-LMMs would provide insights into its adaptability and performance across different architectures.
@@ Appendix B Appendix overview
  • § F More Qualitative Examples.
@@ Appendix C Implementation Details
We conduct Temporal Preference Optimization (TPO) on LongVA [65] and LLaVA-Video [68], two state-of-the-art video-LMMs. The two TPO models are trained using 8 Nvidia A100 80GB GPUs, with a batch size of 64. For preference optimization, we set the KL-divergence weight ($\beta$) to 0.3 and the supervised fine-tuning (SFT) loss weight ($\alpha$) to 0.5 for LongVA-TPO and we set the KL-divergence weight ($\beta$) to 0.2 and the supervised fine-tuning (SFT) loss weight ($\alpha$) to 1 for LLaVA-Video-TPO. We employ full fine-tuning for both the multimodal projector and the language model while keeping the visual encoder frozen, using a learning rate of $4\times 10^{-6}$ for LongVA-TPO and $3\times 10^{-7}$ for LLaVA-Video-TPO. The training is performed on a curated dataset of 10k samples for one epoch for LongVA-TPO and 10k samples for one epoch for LLaVA-Video-TPO. A cosine learning rate scheduler with a warm-up ratio of 0.1 is utilized [38]. The entire TPO fine-tuning process takes approximately 4 hours on both two models.
For evaluation, we adopt the protocol outlined by LongVA [65] and LLaVA-Video [68], leveraging the official lmms-eval repository [64] to assess our model’s performance on three benchmarks. For LongVA-TPO, we set the parameter max_frames_num=128max\_frames\_num=128 across all three benchmarks. For LLaVA-Video-TPO, we set the parameter max_frames_num=96max\_frames\_num=96 for the Video-MME benchmark and max_frames_num=128max\_frames\_num=128 for the rest of the benchmarks.
@@ Appendix D Preference Dataset Curation
We manually curated a set of 200 keywords assisted with GPT-4o-mini [2], which were utilized to retrieve 8,000 videos from the internet, forming a diverse and comprehensive dataset. Using this dataset, we further developed 10,000 queries paired with their corresponding preference responses, covering a broad range of tasks. The detailed prompts for preference dataset curation are provided in Fig. 8 and Fig. 9. For LLaVA-Video, we sampled a subset of 10k QA pairs from the LLaVA-Video-178k dataset with the negative responses only curated by incomplete videos.
The distribution of video lengths in our collected dataset is presented in Fig. 7. The distribution of tasks is illustrated in Fig. 7, encompassing Temporal Reasoning (8.7%), Action Reasoning (12.4%), Causal Reasoning (11.1%), Information Extraction (18.0%), Descriptive Questions (12.8%), Summarization (7.5%), Object Reasoning (14.9%), and Spatial Reasoning (13.5%).
@@ Appendix E Preference Dataset Examples
We provide three additional examples of preference datasets, as illustrated in Fig. 10. For instance, in Example (a), the task involves an OCR-based query aimed at retrieving the quote located beneath a mural. The dis-preferred response incorrectly identifies the relevant frame, failing to locate the quote below the mural and instead referencing another frame containing the phrase “Forward, Warrior.” In contrast, the preferred response accurately identifies the corresponding frame based on the question. This is achieved by leveraging the highly relevant sub-video segment provided to the video-LMM, enabling the correct extraction of both the quote and its attribution.
@@
Appendix F Qualitative Analysis Examples